The probability of an integrated circuit passing a quality test is 0.86. In a batch of 10 integrated circuits find the probabilities that
2 will fail
8 will pass
at least 2 will fail
fewer than 2 will fail.
1
Expert's answer
2021-10-01T15:07:38-0400

Let "X=" the number of integrated circuits which will pass a quality test: "X\sim Bin(n, p)."

Given "n=10, p=0.86, q=1-p=1-0.86=0.14."

a)


"P( \text{2 will fail})=P(X=8)"

"=\dbinom{10}{8}(0.86)^8(0.14)^{10-2}=0.26391"

b)


"P( \text{8 will pass})=P(X=8)"

"=\dbinom{10}{8}(0.86)^8(0.14)^{10-2}=0.26391"



c)


"P( \text{at least 2 will fall})=P(X\leq8)"

"=1-(P(X=9)+P(X=10))"




"=1-\dbinom{10}{9}(0.86)^9(0.14)^{10-9}"

"-\dbinom{10}{10}(0.86)^{10}(0.14)^{10-10}"

"=0.41844"

d)


"P( \text{fewer than 2 will fail})=P(X>8)"

"=P(X=9)+P(X=10)"




"=\dbinom{10}{9}(0.86)^9(0.14)^{10-9}+\dbinom{10}{10}(0.86)^{10}(0.14)^{10-10}"

"=0.58156"